OpenStudy (anonymous):

Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD a month to his collection while Nita is adding 5 CDs a month to her collection. Find the number of months after which they will have the same number of CDs. A. 1 month B. 3 months C. 2 months D. 33 months

OpenStudy (anonymous):

C!? Its right. right? @BacardiDavis

OpenStudy (dan815):

u understand how to build equations right

OpenStudy (anonymous):

ehhh kinda lol

OpenStudy (dan815):

want me to teach u that

OpenStudy (anonymous):

It's D

OpenStudy (anonymous):

yes! that would be aewsome (^.^)

OpenStudy (dan815):

lol trumpter

OpenStudy (anonymous):

I MEAN B MAH BAD

OpenStudy (dan815):

ok lets start

OpenStudy (anonymous):

3 months not 33 read it wrong.

OpenStudy (anonymous):

lol i was lik o.0

OpenStudy (anonymous):

It's 100% 3 months

OpenStudy (anonymous):

thanks @trumpetmaster7777

OpenStudy (dan815):

so we got 1 guy starting with 30 CDs ---> this guys CDS going up 1 every month, so in N months he will have N CDS more and 18 CDS ---> this guys CDS going up 5 every Month, so in N months he will have 5 * N CDS

OpenStudy (anonymous):

np

OpenStudy (dan815):

can u write those 2 equatiosn for me

OpenStudy (dan815):

equation for the first guy and the 2nd guy separetly

OpenStudy (anonymous):

i suck at math (/./)

OpenStudy (dan815):

okay just read those statements then, do the statements make sense

OpenStudy (anonymous):

Yes, like i get it :P

OpenStudy (dan815):

ok, dont worry this is very simple, u dont suck at math trust me :)

OpenStudy (anonymous):

1x + 30 = 5x + 18 =

OpenStudy (dan815):

stop trumptet

OpenStudy (dan815):

he need to understand how to build up his own equation

OpenStudy (anonymous):

alright, sorry :(

OpenStudy (dan815):

lets look at the first guy! lets make an equation for him Let Y = the number of CDs he has Let X = number of months Y = some function of X

OpenStudy (dan815):

read this tell me if its making sense

OpenStudy (anonymous):

y=cds x=months Y???

OpenStudy (dan815):

something not make sense?

OpenStudy (anonymous):

What is the Y for? ohhhh pshhhh CD's??? Y=number of cd's as well?

OpenStudy (dan815):

first we defined what our variables are now, now we need to find the equation

OpenStudy (dan815):

when i say Y is a function of X im plain english iam saying, Y is changing as X is changing, i want to know what Y is for a certain change of X, we need to write an equation for that

OpenStudy (anonymous):

ohh o.0

OpenStudy (dan815):

okay lets start working on equation >_>

OpenStudy (dan815):

look at the initial condition, we know that at the start he has 30 CDS, so even before any time passes he has 30 CDS lets add this to our equation Y=30 now we know when X is changing he get 1 CD for every change in X so for 1 month he has 1 CD more, for 2 months he has 1 * 2 more CDs, for 3 months he has 1*3 CD for X months he has 1*X CD more Y=30+(1*X) Y=30+X one times x is just x

OpenStudy (dan815):

make sense?

OpenStudy (anonymous):

yes! thanx so much!

OpenStudy (anonymous):

about time, lol

OpenStudy (dan815):

okay can you make the equation for the other guy now :)

OpenStudy (anonymous):

lol ikr!

OpenStudy (dan815):

this isnt so simple to get at first, u are doing a very good job understanding :)